OpenStudy (anonymous):

Two integral problems: Explanations and steps please! Integral of (s+1)/(4-s^2)^(1/2) ::I don't know how to deal with the + 1 part. Integral of (tan^-1 x)/(1+x^2) ::I've never dealt with one of these. I tried integration by parts but is stuck with the inverse tangent.

OpenStudy (anonymous):

here are the solution

OpenStudy (anonymous):

1st one you can write as \[\int\limits \frac{s}{\sqrt{4-s^2}}ds+\int\limits \frac{1}{\sqrt{4-s^2}}ds\] Here the 1st one is easy and for the second one use substitution \[s=2\sin(t)\]

OpenStudy (anonymous):

2nd is \[\int\limits (\tan^{-1}x)(1+x)dx\] Use substitution \[t = \tan^{-1}x\]

OpenStudy (anonymous):

sorry i misspelled is tan^(-1)x *(1+x^2)

OpenStudy (anonymous):

second is a set up for u -sub, have \[\int \frac{\tan^{-1}( x)}{1+x^2}dx\] but \[u=\tan(x),du=\frac{dx}{1+x^2}\] and you get \[\int u du\] ready to go

OpenStudy (anonymous):

typo above, should have put \[u=\tan^{-1}(x)\]

OpenStudy (anonymous):

ah , I even wrote the problem wrong :D , Satellite is correct. But the main idea is the substitution \[t = \tan^{-1}x\]

OpenStudy (anonymous):

Oh, thanks. I'm getting stuck in so many exercises.

OpenStudy (anonymous):

Wait, actually I still don't get the second one. If \[u = \tan^-1 (x)\] Won't it become \[\tan^-1 (\tan^-1 x)\] And I don't get how that works because if u is that then \[du = 1/(x^2+1)\] You put it in and it'll become \[\int\limits_{?}^{?}(1+x^2)/(1+x^2)dx\] How does that work? How do pull it out? It's \[\int\limits_{?}^{?} du/(x^2+1)\]

OpenStudy (anonymous):

\[\ u = tan^{-1}x \] \[\ du = \frac{1}{1+x^{2}}dx\] You can now substitute the expression(1/(1+x^2))dx in the integral with du If you don't see it right away, try doing this" \[\ dx = (1+x^2)du\] and now plug this into the integral and you will get" \[\int \frac{u}{1+x^2}(1+x^2)du \]

OpenStudy (anonymous):

Oh, I see. Thank you.